What mass of calcium chloride will be produced from 4.09 g of HCl reacting with Ca(OH)2?
1
Expert's answer
2020-04-24T13:33:44-0400

m(CaCl2) = n(CaCl2) × M(CaCl2) = 0.5n(HCl) × M(CaCl2) = 0.5(m(HCl)/M(HCl)) × M(CaCl2) = 0.5 × 4.09 / 36.46 × 110.98 = 6.22 (g)